how many millimeters of a 40 percent acid solutionn and how many of a 10 percent acid solution must be mixed together to produice 30 millimeters 20 percents solution
Answer by checkley77(12844)

.40X+.10(30-X)=.20*30
.40X+3-.10X=6
.30X=6-3
.30X=3
X=3/.30
X=10 MILLIMETERS OF 40% ACID SOLUTION IS NEEDED.
PROOF:
.40*10+.10(30-10)=.20*30
4+.10*20=6
4+2=6
6=6